A car starts at a speed of 30 meters per second and travels at a uniform acceleration of 10 seconds and achieves a speed of 60 meters per second. How far did the car travel?
1
Expert's answer
2020-12-21T11:00:58-0500

The distance travelled under the uniform acceleration "a" is:


"d =v_it + \dfrac{at^2}{2}"

where "v_i = 30m\/s" is the initial speed, and "t = 10s" is the time.

By definition, acceleration is:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = 60m\/s". Substituting "a" into the first equation, obtain:


"d = v_it + \dfrac12t(v_f - v_i) = \dfrac12t(v_f+v_i)\\
d = \dfrac12\cdot 10\cdot (60+30) = 450m"

Answer. 450 m.